OpenStudy (anonymous):

Solve this problem in log exponential form 3^x=7^x-4

OpenStudy (anonymous):

assuming it is \[3^x=7^{x-4}\] start with \[x\ln(3)=(x-4)\ln(7)\] and then do a raft of algebra to solve for x

OpenStudy (anonymous):

Thank you very much

OpenStudy (mertsj):